OpenStudy (ksaimouli):

solve each system of equations 2x-3y=z -7x+4y=6

OpenStudy (anonymous):

3 variables?

OpenStudy (ksaimouli):

2x-3y=2 -7x+4y=6

OpenStudy (ksaimouli):

how u got

OpenStudy (anonymous):

2x-3y=2 -7x+4y=6 U have to get rid of one of the variables. To do this u can try to make the coefficients of one of the variables the same. If you multiply the first equation by 4 and the second one by 3, then u can add the two new equations together and the y's will disappear.

OpenStudy (ksaimouli):

how we know to multiply 4 and 3

OpenStudy (anonymous):

To make the coefficients of y both equal to 12

OpenStudy (ksaimouli):

can u do please i did but igot x only

OpenStudy (anonymous):

Correct, now u have x put that value into one of your original equations and u will get y.